# Training on test proteins improves fitness, structure, and function prediction

### v1

@@ 1 Introduction
[] Refer to caption
Figure 1: Example of test-time training (TTT) applied to protein folding. ESMFold poorly predicts the structure of the CASP14 target T1074 (shown in white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (Fig. 2E in Lin et al. (2023) and the left panel here). Self-supervised test-time training of ESM2 on the single sequence of T1074 directly minimizes the perplexity, leading to improved structure prediction (better TM-score alignment and higher pLDDT predicted confidence). The same test-time training approach is also broadly applicable to other tasks, such as protein fitness and function prediction.
A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (Subramaniam & Kleywegt, 2022; Tyers & Mann, 2003; Papkou et al., 2023). While machine learning models have demonstrated remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Yang et al., 2024; Kouba et al., 2023). However, biologists often focus their research on individual proteins or protein complexes involved for example in metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signalling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022). In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
Nonetheless, general machine learning models for proteins often struggle to generalize to individual case studies due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Tagasovska et al., 2024; Feng et al., 2024). Bridging the gap between broad, dataset-wide optimizations and the precision required for studying single proteins in practical applications remains a critical challenge in integrating machine learning into biological research (Sapoval et al., 2022).
By contrast, in other application domains of machine learning, such as computer vision and natural language processing, customization and adaptation approaches have emerged as powerful tools to improve model performance in specific contexts (Ruiz et al., 2023; Hardt & Sun, 2023). Drawing inspiration from the test-time training (TTT) approach developed in computer vision to mitigate distribution shifts (Sun et al., 2020; Gandelsman et al., 2022), in this work we propose the TTT approach for proteins. Our method enables adapting protein models to one protein at a time, on the fly, and without the need for additional data. Given a model that has been pre-trained using masked language modeling, our method minimizes the perplexity of the model on a given test protein through self-supervised fine-tuning, which, in turn, results in improved downstream performance without updating the downstream task head.
The prevalence of masked modeling in protein machine learning makes our method broadly applicable to various downstream tasks. Empirically, we demonstrate its effectiveness across three key challenges in protein machine learning. First, TTT achieves state-of-the-art results on the ProteinGym dataset (Notin et al., 2024), a well-established benchmark for protein fitness prediction. Second, TTT enhances protein structure predictions with ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024) on challenging targets. Third, the application of TTT to protein function predictors results in improved classification of terpene synthase (TPS) substrates and protein subcellular localization.
In summary, the key contributions of this work are three-fold:
  1. 1.
    Motivated by the generalization challenges and distribution shifts prevalent in protein machine learning, we introduce a new test-time training (TTT) method [^1] https://github.com/anton-bushuiev/ProteinTTT that enables models to adapt to individual proteins on the fly and without requiring additional data.
  2. 2.
    We establish a link between our TTT approach and perplexity minimization, providing an insight into why this approach enhances model effectiveness.
  3. 3.
    We empirically validate TTT, achieving state-of-the-art results in protein fitness prediction, improving the protein structure prediction capabilities of well-established folding models, and enhancing protein function predictions in the tasks of terpene synthase substrate classification and protein localization prediction.
2 Background and related work
In this section, we present the context and related work that highlight the feasibility and broad applicability of test-time training (TTT) in the domain of machine learning on proteins. The widespread adoption of Y-shaped architectures relying on masked modeling enables the development of a general method for adapting protein models at test time via masking-based self-supervised fine-tuning.
The Y-shaped paradigm of learning.
In machine learning applied to biology, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$, a self-supervised head $g$, and an alternative fine-tuning head $h$. During training, $g\circ f$ is first pre-trained, after which the pre-trained backbone $f$ is reused to fine-tune $h\circ f$ toward a downstream task. Here, $\circ$ denotes a composition of two machine learning modules (e.g., $g$ is applied on top of $f$ in $g\circ f$). At test time, the final model $h\circ f$ is fixed. Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors acquired in the fine-tuning head $h$. This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other protein-related tasks (Hayes et al., 2024).
The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (Hayes et al., 2024). The fine-tuning head $h$, however, depends on the application. In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup, (Meier et al., 2021; Dutton et al., 2024)). In some cases, the fine-tuning head $h$ may also be a machine learning algorithm other than a neural network (Samusevich et al., 2024).
Masked modeling.
While the objective of fine-tuning $h\circ f$ is determined by the downstream application, the choice of pre-training objective for $g\circ f$ is less straightforward. Nevertheless, most methods employ various forms of masked modeling, i.e., optimizing the model weights to accurately reconstruct missing parts of proteins, regardless of the downstream application. Masked modeling pre-traning underpins models for protein structure (Lin et al., 2023) and function (Samusevich et al., 2024) prediction, as well as for protein design (Hayes et al., 2024). For example, in AlphaFold2, most of the loss function weight is put onto masked modeling of multiple sequence alignments (MSAs) (Jumper et al., 2021), and the model has been effectively fine-tuned for various tasks beyond structure prediction (Jing et al., 2024; Cheng et al., 2023; Motmaen et al., 2023).
Masked modeling is a dominant pre-training objective not only across different tasks but also across various protein representations. Sequence models applied to proteins are typically pre-trained to predict randomly masked amino acids in a random or autoregressive manner (Lin et al., 2023; Rao et al., 2021; Elnaggar et al., 2023; Madani et al., 2023; Ferruz et al., 2022; Rives et al., 2021; Rao et al., 2020). Models utilizing graph neural networks or 3D convolutions on protein structures are also commonly pre-trained to fill in missing structural fragments (Dieckhaus et al., 2024; Diaz et al., 2023; Bushuiev et al., 2023; Hsu et al., 2022; Shroff et al., 2020). The most recent approaches combine both sequential and structural information under masked modeling (Hayes et al., 2024; Su et al., 2023; Heinzinger et al., 2023).
Model adaptation.
In many scenarios, machine learning models for proteins benefit from being adapted to a specific protein of interest. This adaptation is commonly achieved in two ways: either via additional input features or via protein-specific fine-tuning. Multiple sequence alignments (MSAs) containing sequences similar to the target protein provide a common way of supplying a model with protein-specific features (Abramson et al., 2024; Jumper et al., 2021; Rao et al., 2021). Another approach for injecting protein-specific knowledge into the model is standard supervised fine-tuning (i.e., via the $h\circ f$ track) on protein-specific data (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019). An alternative is self-supervised fine-tuning (i.e., via the $g\circ f$ track) on proteins from the MSA (Notin et al., 2022b; Frazer et al., 2021; Alley et al., 2019) or on proteins sharing another property with the target protein, such as common family (Sevgen et al., 2023) or class (Samusevich et al., 2024). However, constructing MSAs is time-consuming (Fang et al., 2023), and similar proteins may not be available for many targets (Durairaj et al., 2023; Lin et al., 2023).
Here, we propose an extreme case of self-supervised fine-tuning: learning from a single target protein, without the need for any additional data. To the best of our knowledge, this approach has not been employed in the field of machine learning applied to biology; however, similar methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021). The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is a main inspiration for our work. Here, we demonstrate that TTT is highly relevant for machine learning on proteins even without the presence of explicit distribution shift. We investigate the link of TTT to perplexity minimization and show that TTT improves performance on several downstream tasks.
3 Test-time training (TTT) on proteins
As discussed in the previous section, many machine learning models for proteins employ Y-shaped architectures, consisting of a backbone $f$ with a self-supervised head $g$ and a supervised head $h$. This design facilitates the use of self-supervised fine-tuning across various tasks and models. Notably, most of these models leverage masked modeling as a pre-training objective, which enables the introduction of a broadly applicable test-time training (TTT) method based on masking. Our method adapts models to specific test proteins through masked modeling (Figure 2). In this section, we first formally define the proposed TTT approach (Section 3.1), followed by its application to a range of well-established models (Section 3.2). Finally, we provide an insight into the effectiveness of our method by linking it to perplexity minimization (Section 3.3).
3.1 Self-supervised fine-tuning on test proteins
Figure 2: Overview of our test-time training (TTT) for proteins. Test-time training for proteins builds on the prevalence of Y-shaped architectures relying on masked modeling (i.e., self-supervised masking-based pre-training of $g\circ f$ followed by supervised fine-tuning of $h\circ f$, sharing the backbone $f$). Given a single test protein $x$, TTT adapts the backbone $f$ to the protein using self-supervised fine-tuning. This adaptation leads to better generalization for the downstream task, such as protein fitness, structure, or function prediction.
At test time, we assume a Y-shaped model with a backbone $f$ that has been pre-trained via the self-supervised track $g\circ f$, followed by task-specific fine-tuning through the supervised track $h\circ f$. The goal of test-time training (TTT) is to adapt the backbone $f$ to a single test example $x$ before performing test-time inference on a downstream task via the supervised track.
To achieve this, we first fine-tune the backbone $f$ using the self-supervised track $g\circ f$ on the single example $x$. This step customizes the backbone $f$ to the test sample $x$, and, as demonstrated in Section 4, enhances the generalization of $h\circ f$ without modifying the weights of the task-specific head $h$. Figure 2 illustrates our method. Although the concept of TTT is relatively simple, it involves several important design choices, such as selecting the optimizer and efficiently fine-tuning large backbones, which we describe in the following paragraphs.
Training objective.
We fine-tune $g\circ f$ on a test sample $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
$\mathcal{L}(x)=\mathbb{E}_{M}\Biggl{[}\sum_{i\in M}{-\log{p(x_{i}|x_{\setminus M})}}\Biggl{]},$ (1)
where $x$ denotes a sequence of protein tokens (typically amino acid types), and $\mathbb{E}_{M}$ represents the expectation over randomly sampled masking positions $M$. The loss function $\mathcal{L}(x)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M})}$ of the true tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$. Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be straightforwardly extended to models employing autoregressive masking.
In practice, $\mathbb{E}_{M}$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution) (Hayes et al., 2024). During test-time training, we replicate the masking distribution used during the pre-training. If relevant, we also replicate other pre-training tricks, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).

### v2

# One protein is all you need

@@ 4.1 Protein structure prediction
[…]
Protein structure prediction is the task of predicting 3D coordinates of protein atoms given the amino acid sequence. It is arguably one of the best-established problems in computational biology (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024).
Evaluation setup.
To evaluate the performance of ProteinTTT, we employ CAMEO, a standard benchmark for protein folding. We use the validation and test folds from Lin et al. (2023), focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜E.1). We use the standard TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013) metrics to evaluate global and local structure prediction quality, respectively.
As baseline methods, we use techniques alternative to ProteinTTT for improving the performance of the pre-trained base models. In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023). For each sequence, we sample a number of predictions equal to the total number of ProteinTTT steps and refer to this baseline as ESMFold + MP (Masked Prediction). As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
Results.
Customization with ProteinTTT consistently improves the performance of all the tested methods, ESMFold, HelixFold-Single, and ESM3, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table˜2. Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 5, and 6 structures from ESMFold, HelixFold-Single, and ESM3, respectively, while only slightly disrupting the prediction of 2, 1, and 1 structures, respectively (Figure˜A6). Most notably, ProteinTTT enables accurate structure prediction for targets that are poorly predicted with the original models. For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al. (2023), Fig. 2E). Another example is shown in Figure˜3, where ProteinTTT refined the structure prediction from a low-quality prediction (TM-score =0.29=0.29) to a nearly perfectly folded protein (TM-score =0.92=0.92). Figure˜A4 shows that ESMFold + ProteinTTT maintains computational efficiency of ESMFold, being an order of magnitude faster than AlphaFold2. Figure˜A11 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.
4.2 Protein Fitness Prediction
[…]
The task of protein fitness prediction is to accurately order mutations of a protein based on their disruptive/favorable effects on protein functioning.
Evaluation Setup.
We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (Notin et al., 2024), focusing on its well-established zero-shot setup. Since the zero-shot setup only provides a test set without any data split, we also validate ProteinTTT on independent data. To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019). Following ProteinGym, we measure performance on both datasets using Spearman correlation between predicted and experimental fitness values.
Results.
ProteinTTTconsistently enhances fitness prediction performance of all the tested models across varying model scales (35M and 650M parameters for both ESM2 and SaProt; 110M for ProSST) and both datasets, i.e., test ProteinGym (Table˜4) and validation MaveDB (Table˜A5). Notably, ProSST + ProteinTTT sets a new state of the art on the ProteinGym benchmark (Spearman correlation coefficients calculated for individual deep mutational scanning experiments (DMSs) have statistically significant difference according to a paired t-test with p < 0.05).
We observe that ProteinTTT primarily improves performance for proteins with low MSA depth (i.e., the number of available homologous sequences), suggesting that single-sequence customization enhances predictions for proteins with fewer similar sequences in the training data (Table˜A4). The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (Notin, 2025). We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A5). We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜B).
4.3 Protein function prediction
Finally, we demonstrate a proof of concept for customization in the context of protein function prediction. We experiment with two tasks: predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Samusevich et al., 2025). Appendix˜C shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
5 Case studies
ProteinTTTcan be readily incorporated into structure, fitness, or function prediction pipelines by adding several lines of code (Appendix˜D). Here, we demonstrate two challenging structure prediction case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
5.1 Modeling antibody–antigen loops
Figure 4: ProteinTTT improves modeling of antibody–antigen loops. (a) Average LDDT on the antibody complementarity-determining regions (CDRs, 175 structures) and antigens (814 structures) from the SAbDab dataset with ESMFold pLDDT < 70. Error bars indicate 95% confidence intervals estimated from 1000 bootstrap samples. (b) Example of improved structure prediction for CDRs in the 8K2W entry. The CDR regions H1, H2, and H3, i.e., the parts of the antibody that bind to the antigen, are highlighted with spheres, while black lines show the alignment error between the ground-truth CDR structure (white) and the predictions (colored).
Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (Bennett et al., 2025). However, the presence of highly variable loop regions makes modeling of these interactions a long-standing challenge. Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (Dunbar et al., 2014).
We take the structures from SAbDab that are not predicted well by ESMFold (pLDDT < 70) and show that ProteinTTT improves the LDDT score for 115 of 175 antibody CDR substructures (66%) and 487 of 814 antigen chains (60%). As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold. Figure˜4b illustrates how ProteinTTT enables accurate prediction of all three CDRs in an antibody chain, providing an improved understanding of its binding interface with the corresponding antigen.
5.2 Expanding known structures of viral proteins
Figure 5: ProteinTTT expands the Big Fantastic Virus Database (BFVD). (a) ProteinTTT (light green) substantially improves the performance of ESMFold (yellow) on viral proteins, yielding better structures (pink) for 19% of BFVD entries compared to the original predictions by AlphaFold2 (green). (b) Improvements in pLDDT for ESMFold after ProteinTTT correspond to improvements in LDDT, as benchmarked against BFVD AlphaFold2 structures with pLDDT >> 90. (c) ProteinTTT provides the largest pLDDT improvements (y-axis) for the most out-of-distribution proteins, i.e., those with the smallest MSAs (left on the x-axis) from the Logan database. (d) Structural comparison for BFVD entry UPI000641889E against the PDB structure 2N2J (100% sequence identity) shows that ESMFold + ProteinTTT yields a prediction closest to the ground truth (gray), as also measured by LDDT. (e–g) Additional examples of high-quality viral structures (as measured by pLDDT) predicted with ESMFold + ProteinTTT but not with ESMFold or AlphaFold2. Higher pLDDT values are better.
Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (Bravi, 2024). Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (Kim et al., 2025). Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the repository of viral protein structures (Kim et al., 2025).
Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (Mirdita et al., 2022) using MSAs constructed from Logan (Chikhi et al., 2024), only 55% have high-quality structure predictions (pLDDT > 70). We apply ESMFold and ESMFold + ProteinTTT to the BFVD entries to expand the database with higher-quality structures. This is achieved by applying all three methods to the specific protein and taking the predicted structure with the highest pLDDT. While ESMFold manages to improve the predicted structure (as measured by pLDDT) for 10% of the BFVD proteins, ESMFold + ProteinTTT leads to an improvement for 19% of the dataset entries, substantially increasing the quality of known viral protein structures (Figure˜5a).
We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A9). Notably, the largest improvements in pLDDT align with the largest improvements in LDDT (Figure˜5b). We find that the benefit of customization saturates with the number of homologs available for a protein, indicating that ProteinTTT is most effective for challenging, out-of-distribution proteins (Figure˜5c). Finally, Figure˜5d–g shows examples where ProteinTTT enables high-confidence structure predictions in cases where general-purpose, uncustomized AlphaFold2 and ESMFold struggle.
6 Discussion
We introduce ProteinTTT, a method for customizing protein language models to individual targets. ProteinTTT consistently improves performance across various models, their scales, and downstream tasks. It excels on challenging, out-of-distribution examples where general models often fail. We demonstrate its practical value through two case studies: enhancing the structural prediction of difficult antibody-antigen loops and improving 19% of low-confidence viral protein structures in the Big Fantastic Virus Database. Our work establishes per-protein customization as a powerful and practical tool for biological research.
Acknowledgments
We thank Milot Mirdita for discussions and feedback about this work. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069. This work was also supported by the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560. This work was also supported by the European Union (CLARA (No. 101136607), ERC FRONTIER (no. 101097822), ELIAS (No. 101120237)) and the Technology Agency of the Czech Republic under the NCC Programme from the state budget (No. RETEMED TN02000122) and under the NRP from the EU RRF (No. TEREP TN02000122/001N). This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon Europe program (ERC, TerpenCode, 101170268). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.S. acknowledges support by the National Research Foundation of Korea grants (RS-2020-NR049543, RS-2021-NR061659 and RS-2021-NR056571, RS-2024-00396026), Creative-Pioneering Researchers Program and Novo Nordisk Foundation (NNF24SA0092560).
Appendix
Contents
appendix.Aappendix.Bappendix.Cappendix.Dappendix.Esubsection.E.1subsubsection.E.1.1subsubsection.E.1.2subsubsection.E.1.3subsection.E.2subsubsection.E.2.1subsubsection.E.2.2subsubsection.E.2.3subsection.E.3subsubsection.E.3.1subsubsection.E.3.2subsubsection.E.3.3appendix.Fsubsection.F.1subsection.F.2appendix.Gsubsection.G.1subsection.G.2subsection.G.3
Appendix A Justification of customization via perplexity minimization
Figure A1: Quality of protein structure prediction, as measured by TM-score, correlates with perplexity of the underlying language model on the challenging targets from the CAMEO validation set. Higher TM-scores are associated with lower perplexity, indicating that better predictions are linked to lower uncertainty in the language model’s understanding of the protein sequence.
While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (Liu et al., 2021; Zhao et al., 2023). Here, we offer a potential justification for the effectiveness of ProteinTTT by linking it to perplexity minimization.
Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (Brown, 2020; Chelba et al., 2013). Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (Hayes et al., 2024; Lin et al., 2023). For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
$\text{Perplexity}(x)=\exp\Biggl(\frac{1}{|x|}\sum_{i=1}^{|x|}{-\log{p(x_{i}|x_{\setminus i};\theta)}}\Biggl),$ (3)
where $|x|$ is the length of the input protein sequence $x$ and $p(x_{i}|x_{\setminus i};\theta)$ represents the probability that the model correctly predicts the token $x_{i}$ at position $i$ when it is masked on the input $x_{\setminus i}$. Perplexity ranges from 1 to infinity (the lower, the better), providing an intuitive measure of how well a model fits, on average, tokens in a given sequence. A perplexity value of 1 indicates that the model perfectly fits the sequence, accurately predicting all the true tokens.
Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024). To give an example, we analyze the correlation between perplexity and structure prediction quality (Figure˜A1; see Section˜4.1 for experimental details). A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A10).
Since we assume only a single target example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x;\theta)$ (Equation˜2) on this example is directly linked to minimizing the perplexity Perplexity(x)(x) (Equation˜3). For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity. More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023). As a result, applying self-supervised test-time customization on $x$ through $g\circ f$ enhances the representation of the target protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
Appendix B Customization with multiple sequence alignment (MSA)
Customization training objective.
Since many target proteins may not have homologous sequences (Rao et al., 2021) and finding such homologs may be time-consuming (Lin et al., 2023), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization. However, we also extend the loss function to the case when a multiple sequence alignment (MSA) is available:
$\mathcal{L}_{\text{MSA}}(x;\theta)=\mathbb{E}_{x^{\prime}\sim p_{\text{MSA}}(x^{\prime}|x)}\big[\mathcal{L}(x^{\prime};\theta)\big],$ (4)
[…]
Results for fitness prediction.
Table A1: ProteinTTT can be used with MSA when available. Please see Table˜4 for evaluation details.
Method Avg. Spearman $\uparrow$
ESM2 (Lin et al., 2023)
0.4139
[…]
0.4299 ± 0.00099
MSA Transformer (Rao et al., 2021)
0.4319
MSA Transformer + ProteinTTT (Ours)
0.4326 ± 0.00003
[…]
Appendix C Customization for protein function prediction
Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (Yu et al., 2023; Radivojac & et al., 2013; Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2025). While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (Song et al., 2024), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Christianson, 2017; Samusevich et al., 2025). Using ProteinTTT with EnzymeExplorer (Samusevich et al., 2025) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, we achieve consistent performance gains.
Evaluation setup.
Figure A2: Customization with ProteinTTT enables the correct substrate classification for a terpene synthase (TPS) enzyme. With progressive customization steps of EnzymeExplorer + ProteinTTT, the probability of the initially misclassified substrate (red) decreases, while the probability of the true substrates (green) increases. The bar plots also display the predicted probabilities for other substrates with non-zero values (grey).
For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al. (2025) and reuse the original cross-validation schema. In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from Stärk et al. (2021) as the test set.
Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations. We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (Samusevich et al., 2025): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC). In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
Results.
Table A2: Customization with ProteinTTT improves protein function prediction. For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2025). Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021). The error bars show standard deviations across five random seeds.
TPS substrate classification
Method
mAP $\uparrow$ AUROC $\uparrow$
EnzymeExplorer
(Samusevich et al., 2025)
0.805 0.948
EnzymeExplorer + ProteinTTT
0.811 ± 0.0011 0.950 ± 0.0002
Subcellular localization prediction
Method
Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
Light attention
(Stärk et al., 2021)
0.627 0.549 0.618
Light attention + ProteinTTT
0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
Customization with ProteinTTT improves model performance on both of the protein function prediction tasks and across all considered metrics (Table˜A2). Figure˜A2 provides a qualitative result, where customization with ProteinTTT iteratively refines the prediction of EnzymeExplorer toward a correct TPS substrate class. We hypothesize that improvement with customization is more challenging in classification tasks, as opposed to regression problems, because a larger change in the latent space is required to shift the top-class probability.
Appendix D Implementation details
Infrastructure.
All experiments with ProteinTTT are conducted on machines equipped with a single NVIDIA A100 40GB GPU, an 8-core AMD processor, and 128 GB of physical memory.
Source code.
We provide a user-friendly and easily extensible PyTorch (Paszke, 2019) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT . We provide two Python code snippets ˜1 and ˜2 to demonstrate the implementation of inference and customization with ProteinTTT, respectively. ˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization. Next, ˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class. To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen. The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
Listing 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.
1import esm
2from proteinttt.models.esmfold import ESMFoldTTT, DEFAULT_ESMFOLD_TTT_CFG
3
4# Set protein sequence
5sequence = (
6 "GIHLGELGLLPSTVLAIGYFENLVNIICESLNMLPKLEVSGKEYKKFKFTIVIPKDLDANIKKRAKIY"
7 "FKQKSLIEIEIPTSSRNYPIHIQFDENSTDDILHLYDMPTTIGGIDKAIEMFMRKGHIGKTDQQKLLE"
8 "ERELRNFKTTLENLIATDAFAKEMVEVIIEE"
9)
10
11# Load model
12model = esm.pretrained.esmfold_v1()
13model = model.eval().cuda()
14
15predict_structure(model, sequence)
16# pLDDT: 38.43025
17
18# ============================ ProteinTTT ===============================
19# Customize model to sequence
20model = ESMFoldTTT.ttt_from_pretrained(
21 model, ttt_cfg=DEFAULT_ESMFOLD_TTT_CFG, esmfold_config=model.cfg
22)
23model.ttt(sequence)
24# =======================================================================
25
26predict_structure(model, sequence)
27# pLDDT: 78.69619
28
29# ============================ ProteinTTT ===============================
30# Reset model to original state (after this model.ttt can be called with
31# another protein)
32model.ttt_reset()
33# =======================================================================